### v1

# Differentially Private Distribution Release of Gaussian Mixture Models via KL-Divergence Minimization

Abstract
Gaussian Mixture Models (GMMs) are widely used statistical models for representing multi-modal data distributions, with numerous applications in data mining, pattern recognition, data simulation, and machine learning. However, recent research has shown that releasing GMM parameters poses significant privacy risks, potentially exposing sensitive information about the underlying data. In this paper, we address the challenge of releasing GMM parameters while ensuring differential privacy (DP) guarantees. Specifically, we focus on the privacy protection of mixture weights, component means, and covariance matrices. We propose to use Kullback-Leibler (KL) divergence as a utility metric to assess the accuracy of the released GMM, as it captures the joint impact of noise perturbation on all the model parameters. To achieve privacy, we introduce a DP mechanism that adds carefully calibrated random perturbations to the GMM parameters. Through theoretical analysis, we quantify the effects of privacy budget allocation and perturbation statistics on the DP guarantee, and derive a tractable expression for evaluating KL divergence. We formulate and solve an optimization problem to minimize the KL divergence between the released and original models, subject to a given $(\epsilon,\delta)$-DP constraint. Extensive experiments on both synthetic and real-world datasets demonstrate that our approach achieves strong privacy guarantees while maintaining high utility.
Index Terms:
Gaussian mixture model, density estimation, distribution release, model fitting, differential privacy.
I Introduction
In recent years, the remarkable success of data-driven artificial intelligence (AI) has spurred an increasing demand for the sharing and analysis of large-scale, multi-class, and high-dimensional datasets across a variety of domains, such as healthcare records, consumer transactions, and mobility traces. Organizations have recognized the potential of sharing data statistics to enhance data mining, improve public services, optimize recommendations, and facilitate data simulation [bishop2006pattern]. However, sharing raw data or even their statistics raise significant privacy concerns, especially when sensitive attributes of individuals might be inferred. This tension underscores the need for privacy-preserving mechanisms that allow the release of data statistics without compromising personal or proprietary information [dwork2006calibrating].
Among the various statistical modeling approaches, Gaussian Mixture Models (GMMs) stand out as a versatile tool for representing complex, multivariant, and multi-modal data distributions [figueiredo2002unsupervised]. GMMs are widely applied in data mining, machine learning (ML), and pattern recognition. By modeling the overall distribution as a mixture of several Gaussian components, GMMs naturally capture latent subgroups or clusters in the data, such as different risk profiles in healthcare or distinct spending behaviors in retail. Furthermore, GMMs provide a compact and interpretable parameterization, including component means, covariances, and mixture weights (also known as categorical frequencies), which can be shared more efficiently than raw data records. As a motivating real-world application, GMMs are accurate in representing the statistics of energy consumption data. In particular, the logarithmic values of load demand profiles in Advanced Metering Infrastructure (AMI) data have been empirically shown to fit well to GMMs [9799519]. In the context of energy usage, these load demand profiles represent the amount of electricity consumed over time by a given user or household. When fitted to a GMM, the data naturally forms clusters, with each Gaussian component capturing different usage patterns (e.g., high, medium, and low usage times).
Despite their advantages, releasing GMM parameters directly still poses privacy risks, particularly when the underlying dataset contains sensitive information. An adversary could potentially exploit small mixture components or extreme values to infer specific individuals’ records or their associated class labels. For instance, an adversary could inspect the component means of the GMM model fitted to the AMI dataset and detect changes that could identify the presence of a specific customer. Small perturbations in the model, such as flipping the label of a single customer, could significantly alter the estimated parameters, making it easier for the adversary to infer private information. To mitigate these risks, differential privacy (DP) [dwork2006calibrating, dwork2014algorithmic] offers a robust framework for quantifying and controlling the privacy loss resulting from the inclusion or exclusion of any individual record in a dataset. By introducing carefully calibrated noise, we can estimate mixture model parameters that satisfy strong privacy guarantees while still enabling meaningful statistical analysis.
A central challenge in applying DP mechanisms to GMM fitting is maintaining fidelity between the original non-private data distribution and the released one under a DP constraint. Conventional DP mechanisms typically add artificial noise to perturb the model parameters in order to protect data privacy. However, this perturbation inevitably compromises the accuracy of the released GMM. For instance, the noise added to the Gaussian means could distort the true cluster centers, leading to incorrect clustering and a GMM that fails to represent the underlying data structure. Existing work [7590445] addressed differentially private GMMs by adding white Laplace noise to the mixture weights, means, and covariances individually. This approach models the accuracy of differentially private GMMs using the mean-square errors (MSEs) of individual parameters.
Rather than the MSEs that measure the accuracy of releasing individual parameters, we argue that distribution divergence, specifically Kullback–Leibler (KL) divergence, is another comprehensive metric for quantifying the fidelity of GMMs. KL divergence is a well-established statistical distance metric in information theory that quantifies how much one probability distribution diverges from another [cover1999elements]. When used to characterize the utility of distribution release, KL divergence measures the “information” lost when substituting the true distribution with an approximate, privacy-preserving model. It is well-known that computing maximum likelihood estimation (MLE) is asymptotically equivalent to minimizing the KL divergence between the empirical and true distributions [DL_Goodfellow]. Thus, publishing a data distribution with minimal KL divergence ensures that the released model is as accurate as possible in the maximum-likelihood sense. Moreover, when applied to data simulation and augmentation for ML, KL divergence is closely related to the generalization ability of ML models. Extensive research has established PAC-Bayes bounds for quantifying the generalization error of ML models [MAL-112], which depends critically on the KL divergence between the training and testing data distributions. In this context, KL divergence serves as a powerful metric for evaluating how well an ML model trained on a given data distribution (e.g., a released GMM under privacy constraints) generalizes to the true data, thus providing an important tool for assessing the performance of data synthesis and augmentation in ML training.
KL divergence provides a unified framework for characterizing the combined impact of noise perturbation on the accuracy of the released GMM, taking into account the effects of noise on all model parameters. However, characterizing the KL divergence between the original and DP-preserved distributions, particularly when applying DP mechanisms to the released GMM parameters, remains an open problem. In this work, we show that the KL divergence for GMM parameter release exhibits a closed-form expression, thereby opening new research avenues for designing DP mechanisms that minimize KL divergence while releasing differentially private GMM models.
I-A Our Contributions
Motivated by the above discussions, in this work we propose a differentially private GMM estimation framework that places KL divergence at the core of model evaluation. Specifically, we study the fitting of a multi-class labeled dataset to a GMM whose parameters, including mixture weights, Gaussian means, and covariances, are then released under a given DP requirement. We study the privacy mechanism by adding artificial noise to each parameter following certain controllable statistical distributions. By doing so, we obtain a mechanism design problem that seeks to balance the KL divergence utility and data privacy. We further derive a closed-form expression to evaluate the KL divergence between the released private GMM and the original model. We then present a method to bound the achievable DP level by carefully controlling the noise injection. Through this, we optimize the noise distributions to minimize KL divergence subject to a desired DP constraint. The main contributions of this paper are summarized as follows:
  • $\bullet$
    DP-GMM Model Release Framework: We propose a two-step approach for the differentially private release of GMMs by fitting and releasing GMM model parameters under any given $(\epsilon,\delta)$-DP constraint. Our method first fits the dataset into a GMM and then injects multivariate Gaussian and Wishart noise into the Gaussian means and covariances, respectively. For the discrete-valued mixture weights, we introduce a random mapping mechanism that maps the estimated weights to other feasible mixture weights with a controllable mapping probability. We then formulate the DP mechanism design problem as optimizing the privacy budget and the noise distributions to balance the trade-off between DP and data utility, measured by the KL divergence between the released GMM and the original non-private model.
  • $\bullet$
    Privacy Analysis and Privacy-Utility Trade-Off Characterization: We conduct an $(\epsilon,\delta)$-DP analysis, bounding the combined privacy loss from releasing both the continuous parameters (means and covariances) and the discrete parameters (mixture weights). In addition, we derive a tractable, closed-form expression for the KL divergence between the privatized GMM and the original fitted distribution. This enables us to characterize the trade-off between privacy and the overall fidelity of the released model, providing a clear path to control this balance.
  • $\bullet$
    DP Mechanism Design via Privacy-Constrained KL-Divergence Minimization: Based on our analysis of model utility and privacy, we formulate an optimization problem that allocates privacy budgets and optimizes noise distributions by minimizing the expected KL divergence subject to a given $(\epsilon,\delta)$-DP constraint. While the problem is non-convex, we propose a low-complexity alternating optimization solution to compute a local optimum efficiently.
Through extensive experiments on both synthetic and real-world datasets, we demonstrate how our method preserves model accuracy while satisfying rigorous privacy criteria. Specifically, we show that our approach provides a differentially private GMM that adheres to stringent DP requirements, while maintaining a much lower KL divergence and preserving model fidelity in comparison to existing DP mechanisms.
I-B Organization and Notations
The remainder of this paper is organized as follows. We introduce the system model for differentially private parameter release of GMMs in Section II. In Section LABEL:sec3, we analyze the achievable DP and formulate the DP mechanism design problem. In Section LABEL:sec4, we introduce the proposed solution to the DP mechanism design problem. In Section LABEL:sec5, we present experimental results to evaluate the proposed method. Finally, this paper concludes in Section LABEL:sec6.
Throughout, we use regular, bold small, and bold capital letters to denote scalars, vectors, and matrices, respectively. We use ${\bf X}^{T}$ to denote the transpose of ${\bf X}$, ${\bf X}^{H}$ to denote the conjugate transpose, $\mathop{\rm tr}\nolimits({\bf X})$ to denote the trace, and $|{\bf X}|$ to denote the determinant of ${\bf X}$. We use $x_{i}$ to denote the $i$-th entry of vector ${\bf x}$, $x_{ij}$ or $X_{i,j}$ interchangeably to denote the $(i,j)$-th entry of matrix ${\bf X}$, and ${\bf x}_{j}$ to denote the $j$-th column of ${\bf X}$. The real normal distribution with mean $\boldsymbol{\mu}$ and covariance ${\bf C}$ is denoted by $\mathcal{N}(\boldsymbol{\mu},{\bf C})$, and the cardinality of set $\mathcal{S}$ is denoted by $\lvert\mathcal{S}\rvert$. We use $\lVert\cdot\rVert_{p}$ to denote the $\ell_{p}$ norm, ${\bf I}_{N}$ to denote the $N\times N$ identity matrix, $\bf 1$ (or $\bf 0$) to denote the all-one (or all-zero) vector with an appropriate size. We use $x\sim p(x)$ to represent that the random variable $x$ is drawn from the distribution $p(x)$.
II Differentially Private Distribution Release for GMMs
Consider a dataset $\mathcal{D}=\{{\bf x}_{n},y_{n}\}_{n=1}^{N}$, consisting of $N$ labeled samples ${\bf x}_{1},\dots,{\bf x}_{N}\in\mathbb{R}^{d}$ partitioned into $K$ classes. Each sample ${\bf x}_{n}$ is associated with a class label $y_{n}\in\{1,\dots,K\}$. [^1] When unavailable, the class labels can be calculated a priori by using clustering methods such as $K$-Means clustering. Without loss of generality, we assume every class has at least one data point. Our aim is to publicly release a distributional model for these multi-class data in a differentially private manner by fitting them with a multivariate GMM.
To achieve this, we estimate the parameters of the following GMM based on $\mathcal{D}$:
$p(y=k)=\pi_{k}(\mathcal{D}),p({\bf x}|y=k)=\mathcal{N}\left(\boldsymbol{\mu}_{k}(\mathcal{D}),\boldsymbol{\Sigma}_{k}(\mathcal{D})\right),\forall k,$ (1)
where $\pi_{k}(\mathcal{D})>0$ denotes the empirical frequency (mixture weight) of class $k$ satisfying $\sum_{k}\pi_{k}(\mathcal{D})=1$, and $\boldsymbol{\mu}_{k}(\mathcal{D})$ and $\boldsymbol{\Sigma}_{k}(\mathcal{D})$ represent the mean and covariance of the $k$-th Gaussian component, respectively. For brevity, when the dataset $\mathcal{D}$ is implied by context, we omit the argument in the notation.
To satisfy a prescribed DP constraint, we employ a two-step strategy, by first estimating the parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ from the dataset $\mathcal{D}$, followed by applying a DP mechanism before releasing these parameters. Concretely, we compute the parameters via histograms and sample statistics as follows:
$\pi_{k}=\frac{1}{N}\sum_{y_{n}=k}1,$ (2a)
$\boldsymbol{\mu}_{k}=\frac{1}{N\pi_{k}}\sum_{y_{n}=k}{\bf x}_{n},$ (2b)
$\boldsymbol{\Sigma}_{k}=\frac{1}{N\pi_{k}-1}\sum_{y_{n}=k}({\bf x}_{n}-\boldsymbol{\mu}_{k})({\bf x}_{n}-\boldsymbol{\mu}_{k})^{T}.$ (2c)
Define $\boldsymbol{\pi}=[\pi_{1},\dots,\pi_{K}]$. From (2a), it follows that $\boldsymbol{\pi}$ is a discrete histogram belonging to the following set:
$\mathcal{S}=\left\{\boldsymbol{\pi}:\sum_{k=1}^{K}\pi_{k}=1,\pi_{k}\in\left\{\frac{1}{N},\frac{2}{N},\cdots,\frac{N-1}{N}\right\},\forall k\right\}.$ (3)
By the stars-and-bars theorem, the cardinality of $\mathcal{S}$ is given by $|\mathcal{S}|={{N-1}\choose{K-1}}$. Unless otherwise specified, when we referred to an element $\boldsymbol{\pi}\in\mathcal{S}$, we assume a fixed ordering of the elements in $\mathcal{S}$ and thus any feasible $\boldsymbol{\pi}$ has a unique index in the order.
After estimating the parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$, we apply a post-processing step to satisfy given DP constraints before making these parameters public. The following definition of DP, introduced in [dwork2006calibrating, dwork2014algorithmic], is standard in the literature.
Definition 1 ($(\epsilon,\delta)$-DP).
Consider a randomized mechanism $M$ that takes a dataset $\mathcal{D}$ as input and outputs a query answer $M(\mathcal{D})$. Let $\epsilon>0$ and $\delta\in[0,1]$ be given parameters. $M$ is said to satisfy $(\epsilon,\delta)$-DP if, for any two adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ where $\mathcal{D}^{\prime}$ differs from $\mathcal{D}$ by altering the class label of exactly one data point [^2] For simplicity, we define DP in terms of an adjacent dataset $\mathcal{D}^{\prime}$ that alters exactly one class label from $\mathcal{D}$. The same principle extends to the variant where $\mathcal{D}^{\prime}$ replaces a single data point in $\mathcal{D}$ with a bounded alternative or removes/adds one data point from $\mathcal{D}$., the following inequality holds for all measurable sets $\mathcal{H}$:
$\mathrm{Pr}(M(\mathcal{D})\in\mathcal{H})\leq e^{\epsilon}\mathrm{Pr}(M(\mathcal{D}^{\prime})\in\mathcal{H})+\delta,\forall\mathcal{H}.$ (4)
In this work, $M$ refers to the mechanism that releases the GMM parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. Intuitively, an $(\epsilon,\delta)$-DP mechanism makes it difficult for any potential attacker to detect the class label of any single data point in $\mathcal{D}$, especially when $\epsilon$ and $\delta$ are small. Here, $\epsilon$ controls the overall privacy stringency, while $\delta$ governs the probability that this privacy guarantee may not hold.
To satisfy a given DP requirement, we leverage well-known DP mechanisms that add artificial noise to $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ prior to release. Unlike canonical mechanisms designed for unbounded scalars, we must ensure that the added noise respects the dimensionality and retains the feasibility of these parameters. Concretely, we need the perturbed frequency parameters to remain in $\mathcal{S}$, and the perturbed covariance matrices to stay positive semidefinite (PSD).
Toward this end, we assign each parameter a specific noise distribution. For the continuous-valued mean and covariance of each Gaussian component, we adopt a multivariate Gaussian mechanism and a Wishart mechanism [Jiang_Xie_Zhang_2016]. The perturbed means and covariances are given by
$\widetilde{\boldsymbol{\mu}}_{k}=\boldsymbol{\mu}_{k}+{\bf w}_{k},{\bf w}_{k}\sim\mathcal{N}({\bf 0},\boldsymbol{\Gamma}_{k}^{-1}),\forall k,$ (5)
$\widetilde{\boldsymbol{\Sigma}}_{k}=\boldsymbol{\Sigma}_{k}+{\bf W}_{k},{\bf W}_{k}\sim\mathcal{W}_{d}(\frac{1}{\gamma_{k}}{\bf I}_{d},d+1),\forall k,$ (6)
where ${\bf w}_{k}$ and ${\bf W}_{k}$ are the noise terms added to the mean and covariance of the $k$-th component, respectively, $\boldsymbol{\Gamma}_{k}$ is the precision matrix of the Gaussian noise, $\mathcal{W}_{d}(\gamma_{k}{\bf I}_{d},d+1)$ is a Wishart distribution over $d\times d$ matrices with $d+1$ degrees of freedom and scale matrix $\gamma_{k}^{-1}{\bf I}_{d}$, and the parameter $\gamma_{k}>0$ controls the variance of the Wishart noise. Because Wishart random matrices are always positive definite, the construction in (6) ensures that the perturbed covariance remains feasible.
Meanwhile, to privatize the discrete-valued frequency vector $\boldsymbol{\pi}$, we release a random vector $\widetilde{\boldsymbol{\pi}}\in\mathcal{S}$ according to a tunable conditional probability mass function (PMF) $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$. With a fixed ordering of elements in $\mathcal{S}$, this PMF can be viewed as a transition matrix ${\bf F}\in[0,1]^{|\mathcal{S}|\times|\mathcal{S}|}$ whose entry $F_{i,j}$ specifies the probability of mapping the $j$-th input element in $\mathcal{S}$ to the $i$-th output element in $\mathcal{S}$. In the sequel, we will use $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$ and ${\bf F}$ interchangeably to denote the mapping PMF.