### v2

# Differentially Private Distribution Release of Gaussian Mixture Models via KL-Divergence Minimization

Abstract
Gaussian Mixture Models (GMMs) are widely used statistical models for representing multi-modal data distributions, with numerous applications in data mining, pattern recognition, data simulation, and machine learning. However, recent research has shown that releasing GMM parameters poses significant privacy risks, potentially exposing sensitive information about the underlying data. In this paper, we address the challenge of releasing GMM parameters while ensuring differential privacy (DP) guarantees. Specifically, we focus on the privacy protection of mixture weights, component means, and covariance matrices. We propose to use Kullback-Leibler (KL) divergence as a utility metric to assess the accuracy of the released GMM, as it captures the joint impact of noise perturbation on all the model parameters. To achieve privacy, we introduce a DP mechanism that adds carefully calibrated random perturbations to the GMM parameters. Through theoretical analysis, we quantify the effects of privacy budget allocation and perturbation statistics on the DP guarantee, and derive a tractable expression for evaluating KL divergence. We formulate and solve an optimization problem to minimize the KL divergence between the released and original models, subject to a given $(\epsilon,\delta)$-DP constraint. Extensive experiments on both synthetic and real-world datasets demonstrate that our approach achieves strong privacy guarantees while maintaining high utility.
I Introduction
In recent years, the remarkable success of data-driven artificial intelligence (AI) has spurred an increasing demand for the sharing and analysis of large-scale, multi-class, and high-dimensional datasets across a variety of domains, such as healthcare records, consumer transactions, and mobility traces. Organizations have recognized the potential of sharing data statistics to enhance data mining, improve public services, optimize recommendations, and facilitate data simulation [1]. However, sharing raw data or even their statistics raise significant privacy concerns, especially when sensitive attributes of individuals might be inferred. This tension underscores the need for privacy-preserving mechanisms that allow the release of data statistics without compromising personal or proprietary information [2].
Among the various statistical modeling approaches, Gaussian Mixture Models (GMMs) stand out as a versatile tool for representing complex, multivariant, and multi-modal data distributions [3]. GMMs are widely applied in data mining, machine learning (ML), and pattern recognition. By modeling the overall distribution as a mixture of several Gaussian components, GMMs naturally capture latent subgroups or clusters in the data, such as different risk profiles in healthcare or distinct spending behaviors in retail. Furthermore, GMMs provide a compact and interpretable parameterization, including component means, covariances, and mixture weights (also known as categorical frequencies), which can be shared more efficiently than raw data records. As a motivating real-world application, GMMs are accurate in representing the statistics of energy consumption data. In particular, the logarithmic values of load demand profiles in Advanced Metering Infrastructure (AMI) data have been empirically shown to fit well to GMMs [4]. In the context of energy usage, these load demand profiles represent the amount of electricity consumed over time by a given user or household. When fitted to a GMM, the data naturally forms clusters, with each Gaussian component capturing different usage patterns (e.g., high, medium, and low usage times).
Despite their advantages, releasing GMM parameters directly still poses privacy risks, particularly when the underlying dataset contains sensitive information. An adversary could potentially exploit small mixture components or extreme values to infer specific individuals’ records or their associated class labels. For instance, an adversary could inspect the component means of the GMM model fitted to the AMI dataset and detect changes that could identify the presence of a specific customer. Small perturbations in the model, such as flipping the label of a single customer, could significantly alter the estimated parameters, making it easier for the adversary to infer private information. To mitigate these risks, differential privacy (DP) [2, 5] offers a robust framework for quantifying and controlling the privacy loss resulting from the inclusion or exclusion of any individual record in a dataset. By introducing carefully calibrated noise, we can estimate mixture model parameters that satisfy strong privacy guarantees while still enabling meaningful statistical analysis.
A central challenge in applying DP mechanisms to GMM fitting is maintaining fidelity between the original non-private data distribution and the released one under a DP constraint. Conventional DP mechanisms typically add artificial noise to perturb the model parameters in order to protect data privacy. However, this perturbation inevitably compromises the accuracy of the released GMM. For instance, the noise added to the Gaussian means could distort the true cluster centers, leading to incorrect clustering and a GMM that fails to represent the underlying data structure. Existing work [6] addressed differentially private GMMs by adding white Laplace noise to the mixture weights, means, and covariances individually. This approach models the accuracy of differentially private GMMs using the mean-squared errors (MSEs) of individual parameters.
Rather than the MSEs that measure the accuracy of releasing individual parameters, we argue that distribution divergence, specifically Kullback–Leibler (KL) divergence, is another comprehensive metric for quantifying the fidelity of GMMs. KL divergence is a well-established statistical distance metric in information theory that quantifies how much one probability distribution diverges from another [7]. When used to characterize the utility of distribution release, KL divergence measures the “information” lost when substituting the true distribution with an approximate, privacy-preserving model. It is well-known that computing maximum likelihood estimation (MLE) is asymptotically equivalent to minimizing the KL divergence between the empirical and true distributions [8]. Thus, publishing a data distribution with minimal KL divergence ensures that the released model is as accurate as possible in the maximum-likelihood sense. Moreover, when applied to data simulation and augmentation for ML, KL divergence is closely related to the generalization ability of ML models. Extensive research has established PAC-Bayes bounds for quantifying the generalization error of ML models [9], which depends critically on the KL divergence between the training and testing data distributions. In this context, KL divergence serves as a powerful metric for evaluating how well an ML model trained on a given data distribution (e.g., a released GMM under privacy constraints) generalizes to the true data, thus providing an important tool for assessing the performance of data synthesis and augmentation in ML training.
KL divergence provides a unified framework for characterizing the combined impact of noise perturbation on the accuracy of the released GMM, taking into account the effects of noise on all model parameters. However, characterizing the KL divergence between the original and DP-preserved distributions, particularly when applying DP mechanisms to the released GMM parameters, remains an open problem. In this work, we show that the KL divergence for GMM parameter release exhibits a closed-form expression under specific DP mechanisms, thereby opening new research avenues for designing DP mechanisms that minimize KL divergence while releasing differentially private GMM models.
I-A Our Contributions
Motivated by the above discussions, in this work we propose a differentially private GMM estimation framework that places KL divergence at the core of model evaluation. Specifically, we study the fitting of a multi-class labeled dataset to a GMM whose parameters, including mixture weights, Gaussian means, and covariances, are then released under a given DP requirement. We study the privacy mechanism by adding artificial noise to each parameter following certain controllable statistical distributions. By doing so, we obtain a mechanism design problem that seeks to balance the KL divergence utility and data privacy. We further derive a closed-form expression to evaluate the KL divergence between the released private GMM and the original model. We then present a method to bound the achievable DP level by carefully controlling the noise injection. Through this, we optimize the noise distributions to minimize KL divergence subject to a desired DP constraint. The main contributions of this paper are summarized as follows:
  • $\bullet$
    DP-GMM Model Release Framework: We propose a two-step approach for the differentially private release of GMMs by fitting and releasing GMM model parameters under any given $(\epsilon,\delta)$-DP constraint. Our method first fits the dataset into a GMM and then injects multivariate Gaussian and Wishart noise into the Gaussian means and covariances, respectively. For the discrete-valued mixture weights, we introduce a random mapping mechanism that maps the estimated weights to other feasible mixture weights with a controllable mapping probability. We then formulate the DP mechanism design problem as optimizing the privacy budget and the noise distributions to balance the trade-off between DP and data utility, measured by the KL divergence between the released GMM and the original non-private model.
  • $\bullet$
    Privacy Analysis and Privacy-Utility Trade-Off Characterization: We conduct an $(\epsilon,\delta)$-DP analysis, bounding the combined privacy loss from releasing both the continuous parameters (means and covariances) and the discrete parameters (mixture weights). In addition, we derive a tractable, closed-form expression for the KL divergence between the privatized GMM and the original fitted distribution. This enables us to characterize the trade-off between privacy and the overall fidelity of the released model, providing a clear path to control this balance.
  • $\bullet$
    DP Mechanism Design via Privacy-Constrained KL-divergence Minimization: Based on our analysis of model utility and privacy, we formulate an optimization problem that allocates privacy budgets and optimizes noise distributions by minimizing the expected KL divergence subject to a given $(\epsilon,\delta)$-DP constraint. Crucially, this formulation makes the privacy–utility trade-off transparent: the privacy budget is reallocated across DP randomization parameters during the alternating updates, directly exposing how noise levels and budget shares affect the utility. While the problem is non-convex, we propose a low-complexity alternating optimization solution to compute a local optimum efficiently.
Through extensive experiments on both synthetic and real-world datasets, we demonstrate how our method preserves model accuracy while satisfying rigorous privacy criteria. Specifically, we show that our approach provides a differentially private GMM that adheres to stringent DP requirements, while maintaining a much lower KL divergence and preserving model fidelity in comparison to existing DP mechanisms.
I-B Organization and Notations
The remainder of this paper is organized as follows. We introduce the system model for differentially private parameter release of GMMs in Section II. In Section III, we analyze the achievable DP and formulate the DP mechanism design problem. In Section IV, we introduce the proposed solution to the DP mechanism design problem. In Section V, we present experimental results to evaluate the proposed method. Finally, this paper concludes in Section VI.
Throughout, we use regular, bold small, and bold capital letters to denote scalars, vectors, and matrices, respectively. We use ${\bf X}^{T}$ to denote the transpose of $\bf X$}, ${\bf X}^{H}$ to denote the conjugate transpose, $\mathop{\rm tr}\nolimits({\bf X})$ to denote the trace, and $|{\bf X}|$ to denote the determinant of $\bf X$}. We use $x_{i}$ to denote the $i$-th entry of vector $\bf x$}, $x_{ij}$ or $X_{i,j}$ interchangeably to denote the $(i,j)$-th entry of matrix $\bf X$}, and ${\bf x}_{j}$ to denote the $j$-th column of $\bf X$}. The real normal distribution with mean $\boldsymbol{\mu}$ and covariance $\bf C$} is denoted by $\mathcal{N}(\boldsymbol{\mu},{\bf C})$, and the cardinality of set $\mathcal{S}$ is denoted by $\lvert\mathcal{S}\rvert$. We use $\lVert\cdot\rVert_{p}$ to denote the $\ell_{p}$ norm, ${\bf I}_{N}$ to denote the $N\times N$ identity matrix, $\bf 1$ (or $\bf 0$) to denote the all-one (or all-zero) vector with an appropriate size. We use $x\sim p(x)$ to represent that the random variable $x$ is drawn from the distribution $p(x)$.
II Differentially Private Distribution Release for GMMs
Consider a dataset $\mathcal{D}=\{{\bf x}_{n},y_{n}\}_{n=1}^{N}$, consisting of $N$ labeled samples ${\bf x}_{1},\dots,{\bf x}_{N}\in\mathbb{R}^{d}$ partitioned into $K$ classes. Each sample ${\bf x}_{n}$ is associated with a class label $y_{n}\in\{1,\dots,K\}$. [^1] When unavailable, the class labels can be calculated a priori by using clustering methods such as $K$-Means clustering. Without loss of generality, we assume every class has at least one data point. Our aim is to publicly release a distributional model for these multi-class data in a differentially private manner by fitting them with a multivariate GMM.
To achieve this, we estimate the parameters of the following GMM based on $\mathcal{D}$:
$p(y=k)=\pi_{k}(\mathcal{D}),p({\bf x}|y=k)=\mathcal{N}\left(\boldsymbol{\mu}_{k}(\mathcal{D}),\boldsymbol{\Sigma}_{k}(\mathcal{D})\right),\forall k,$ (1)
where $\pi_{k}(\mathcal{D})>0$ denotes the empirical frequency (mixture weight) of class $k$ satisfying $\sum_{k}\pi_{k}(\mathcal{D})=1$, and $\boldsymbol{\mu}_{k}(\mathcal{D})$ and $\boldsymbol{\Sigma}_{k}(\mathcal{D})$ represent the mean and covariance of the $k$-th Gaussian component, respectively. For brevity, when the dataset $\mathcal{D}$ is implied by context, we omit the argument in the notation.
To satisfy a prescribed DP constraint, we employ a two-step strategy, by first estimating the parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ from the dataset $\mathcal{D}$, followed by applying a DP mechanism before releasing these parameters. Concretely, we compute the parameters via histograms and sample statistics as follows:
$\pi_{k}=\frac{1}{N}\sum_{y_{n}=k}1,$ (2a)
$\boldsymbol{\mu}_{k}=\frac{1}{N\pi_{k}}\sum_{y_{n}=k}{\bf x}_{n},$ (2b)
$\boldsymbol{\Sigma}_{k}=\frac{1}{N\pi_{k}-1}\sum_{y_{n}=k}({\bf x}_{n}-\boldsymbol{\mu}_{k})({\bf x}_{n}-\boldsymbol{\mu}_{k})^{T}.$ (2c)
Define $\boldsymbol{\pi}=[\pi_{1},\dots,\pi_{K}]$. From (2a), it follows that $\boldsymbol{\pi}$ is a discrete histogram belonging to the following set:
$\mathcal{S}=\left\{\boldsymbol{\pi}:\sum_{k=1}^{K}\pi_{k}=1,\pi_{k}\in\left\{\frac{1}{N},\frac{2}{N},\cdots,\frac{N-1}{N}\right\},\forall k\right\}.$ (3)
By the stars-and-bars theorem, the cardinality of $\mathcal{S}$ is given by $|\mathcal{S}|={{N-1}\choose{K-1}}$. Unless otherwise specified, when we referred to an element $\boldsymbol{\pi}\in\mathcal{S}$, we assume a fixed ordering of the elements in $\mathcal{S}$ and thus any feasible $\boldsymbol{\pi}$ has a unique index in the order.
After estimating the parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$, we apply a post-processing step to satisfy given DP constraints before making these parameters public. The following definition of DP, introduced in [2, 5], is standard in the literature.
Definition 1 ($(\epsilon,\delta)$-DP).
Consider a randomized mechanism $M$ that takes a dataset $\mathcal{D}$ as input and outputs a query answer $M(\mathcal{D})$. Let $\epsilon>0$ and $\delta\in[0,1]$ be given parameters. $M$ is said to satisfy $(\epsilon,\delta)$-DP if, for any two adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ where $\mathcal{D}^{\prime}$ differs from $\mathcal{D}$ by altering the class label of exactly one data point, the following inequality holds for all measurable sets $\mathcal{H}$:
$\mathrm{Pr}(M(\mathcal{D})\in\mathcal{H})\leq e^{\epsilon}\mathrm{Pr}(M(\mathcal{D}^{\prime})\in\mathcal{H})+\delta,\forall\mathcal{H}.$ (4)
In this work, $M$ refers to the mechanism that releases the GMM parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. Intuitively, an $(\epsilon,\delta)$-DP mechanism makes it difficult for any potential attacker to detect the class label of any single data point in $\mathcal{D}$, especially when $\epsilon$ and $\delta$ are small. Here, $\epsilon$ controls the overall privacy stringency, while $\delta$ governs the probability that this privacy guarantee may not hold.
Remark 1.
Definition 1 adopts label-level DP, where an adjacent dataset $\mathcal{D}^{\prime}$ alters exactly one class label. Record-level DP is also possible under bounded features, with adjacency defined by adding, removing, or replacing one data point. Section IV-C shows how our approach adapts to record-level DP with the same design principles.
To satisfy a given DP requirement, we leverage well-known DP mechanisms that add artificial noise to $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ prior to release. Unlike canonical mechanisms designed for unbounded scalars, we must ensure that the added noise respects the dimensionality and retains the feasibility of these parameters. Concretely, we need the perturbed frequency parameters to remain in $\mathcal{S}$, and the perturbed covariance matrices to stay positive semidefinite (PSD).
Toward this end, we assign each parameter a specific noise distribution. For the continuous-valued mean and covariance of each Gaussian component, we adopt a multivariate Gaussian mechanism and a Wishart mechanism [10]. The perturbed means and covariances are given by
$\widetilde{\boldsymbol{\mu}}_{k}=\boldsymbol{\mu}_{k}+{\bf w}_{k},{\bf w}_{k}\sim\mathcal{N}({\bf 0},\boldsymbol{\Gamma}_{k}^{-1}),\forall k,$ (5)
$\widetilde{\boldsymbol{\Sigma}}_{k}=\boldsymbol{\Sigma}_{k}+{\bf W}_{k},{\bf W}_{k}\sim\mathcal{W}_{d}(\frac{1}{\gamma_{k}}{\bf I}_{d},d+1),\forall k,$ (6)
where ${\bf w}_{k}$ and ${\bf W}_{k}$ are the noise terms added to the mean and covariance of the $k$-th component, respectively, $\boldsymbol{\Gamma}_{k}$ is the precision matrix of the Gaussian noise, $\mathcal{W}_{d}(\gamma_{k}^{-1}{\bf I}_{d},d+1)$ is a Wishart distribution over $d\times d$ matrices with $d+1$ degrees of freedom and scale matrix $\gamma_{k}^{-1}{\bf I}_{d}$, and the parameter $\gamma_{k}>0$ controls the variance of the Wishart noise. Because Wishart random matrices are always positive definite, the construction in (6) ensures that the perturbed covariance remains feasible.
Meanwhile, to privatize the discrete-valued frequency vector $\boldsymbol{\pi}$, we release a random vector $\widetilde{\boldsymbol{\pi}}\in\mathcal{S}$ according to a tunable conditional probability mass function (PMF) $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$. With a fixed ordering of elements in $\mathcal{S}$, this PMF can be viewed as a transition matrix ${\bf F}\in[0,1]^{|\mathcal{S}|\times|\mathcal{S}|}$ whose entry $F_{i,j}$ specifies the probability of mapping the $j$-th input element in $\mathcal{S}$ to the $i$-th output element in $\mathcal{S}$. In the sequel, we will use $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$ and $\bf F$} interchangeably to denote the mapping PMF.
Remark 2.
Our randomization mechanism for the frequency parameter $\widetilde{\boldsymbol{\pi}}$ presumes a discrete-valued input $\boldsymbol{\pi}$, matching the histogram-based parameter estimation in (2a). Other GMM estimation methods, such as the expectation-maximization algorithm [11] may yield continuous-valued frequencies, which can be discretized beforehand if needed.
Our goal is to choose $\{\boldsymbol{\Gamma}_{k},\gamma_{k}\}_{k=1}^{K}$ and $\bf F$} so that the released parameters $\{\tilde{\pi}_{k},\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$ satisfy the $(\epsilon,\delta)$-DP. An end user can then reconstruct the private GMM as
$\widetilde{p}(y=k)=\tilde{\pi}_{k},\widetilde{p}({\bf x}|y=k)=\mathcal{N}(\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}),\forall k.$ (7)
We use the expected KL divergence, to measure the fidelity of $\widetilde{p}({\bf x},y)$:
$\mathbb{E}\left[KL\left(\widetilde{p}({\bf x},y)||p({\bf x},y)\right)\right]=\mathbb{E}\left[\int\widetilde{p}({\bf x},y)\ln\Big(\frac{\widetilde{p}({\bf x},y)}{p({\bf x},y)}\Big)d{\bf x}dy\right],$ (8)
where the expectations are taken with respect to the randomness in $\{{\bf w}_{k},{\bf W}_{k}\}_{\forall k}$ and $\widetilde{\boldsymbol{\pi}}$. A smaller KL divergence value implies that $\widetilde{p}({\bf x})$ is closer to $p({\bf x})$ and therefore a more accurate estimate of the data density.
The following examples demonstrate the effectiveness of the KL divergence for measuring the accuracy of distribution estimation and release.
Example 1 (Classification).
Classification is a fundamental problem in data mining and ML. By fitting a GMM to the raw data, one can capture latent structures or subgroups and use the model to classify new data. However, if the private GMM diverges significantly from the true distribution, these classifiers may be distorted or merged incorrectly. For instance, consider a hospital that privately releases classification information about patients based on clinical measurements (high/low blood pressure, high/low cholesterol, etc.). Policymakers rely on accurate classification boundaries to identify high-risk subpopulations and allocate resources effectively. A high KL divergence between the true and released distributions would imply that the separations are significantly altered, undermining the utility of the published model. Hence, minimizing the KL divergence in (8) ensures that the core data structure is well preserved despite the noise added for privacy.
Example 2 (Density Estimation and Data Simulations).
GMMs are widely used to estimate data density and generate synthetic samples that mimic the statistical properties of the real data. KL divergence is a standard metric for accuracy assessment and model selection of density estimation. Consider a public health agency creating a synthetic version of a patient dataset for open research. If the released GMM has a distribution that significantly diverges from the true one in terms of the KL divergence, the synthetic data may incorrectly represent disease prevalence or demographic proportions. This leads to flawed insights, misleading model development, and potential misallocation of medical resources. By monitoring and minimizing the KL divergence, the agency ensures the synthetic dataset remains faithful to genuine population patterns, while still respecting strict privacy requirements.
Example 3 (Data Augmentation for ML).
Data augmentation plays a critical role in preventing model overfitting and enhancing model generalization, especially in applications where labeled data are expensive to obtain. A differentially private GMM allows organizations to share model parameters that can be used for sampling additional training data while preserving individual privacy. In machine learning theory, PAC-Bayes methods are commonly used to quantify generalization ability when training and testing data follow different distributions. Typical PAC-Bayes bounds indicate that the model generalization error is bounded by a non-diminishing term related to the KL divergence between the training and testing data distributions [9].
Moreover, in the context of data augmentation or synthesis, [12] reported that the generalization error of a model trained with an augmented or synthetic dataset is upper-bounded by a term determined by the total variation distance (TVD) between the true data distribution and the augmented data distribution. However, this TVD-based bound is often intractable, even for simple distribution models such as GMMs. According to Pinsker’s inequality [13], the TVD between two distributions is tightly upper-bounded by the square root of half their KL divergence. In other words, KL divergence serves as a surrogate for TVD and provides a more tractable expression for bounding the model generalization ability. Maintaining low KL divergence ensures that the augmented dataset closely aligns with the true data distribution, resulting in stronger model generalization and more reliable predictive outcomes.
Scope (model-release privacy). Our goal is to protect model-release privacy rather than sample-level prediction privacy. We assume a trusted data owner fits the GMM on the raw dataset and then releases differentially private parameters $\{\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k},\boldsymbol{\pi}\}$ to an untrusted query recipient. The DP guarantees we provide apply to the act of releasing these parameters. Addressing per-query or per-sample prediction privacy is orthogonal to this work and left for future study.
III Problem Formulation
In this section, we present a tractable problem formulation for DP mechanism design by minimizing the KL divergence subject to DP constraints.
III-A DP Analysis
We begin by examining the conditions under which releasing $\{\tilde{\pi}_{k},\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$ achieves a given $(\epsilon,\delta)$-DP requirement, as a function of the controllable parameters $\{\boldsymbol{\Gamma}_{k},\gamma_{k}\}_{\forall k}$ and $f(\widetilde{\boldsymbol{\pi}}\mid\boldsymbol{\pi})$. The core idea is to bound the overall privacy loss by composing the individual privacy contributions of each parameter. Following the analysis in [4, 10], we derive a sufficient condition for satisfying (4) when releasing $\{\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$. We then enumerate all possible adjacent datasets to drive the expression of (4) for $\widetilde{\boldsymbol{\pi}}$.
Given the dataset $\mathcal{D}$ of size $N$, let $\mathcal{D}^{\prime}_{n,k^{\prime}}$ denote the $(n,k^{\prime})$-th adjacent dataset obtained by changing the class label of the $n$-th data point to $k^{\prime}$. There are N(K1)N(K-1) such adjacent datasets in total. A sufficient condition for satisfying $(\epsilon,\delta)$-DP is summarized below.
Theorem 1.
Consider the non-private parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ defined in (2). Releasing the perturbed parameters $\{\tilde{\pi}_{k},\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$ in (5) and (6) satisfies $(\epsilon,\delta)$-DP if $\epsilon$ and $\delta$ fulfill the following inequalities:
$\epsilon\geq\max_{k=1}^{K}\left\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\right\}+\epsilon_{0},$ (9)
$\frac{\epsilon_{k}^{2}}{2\ln(2/\delta)}$
$\penalty 10000\ \penalty 10000\ \geq\sup_{n,k^{\prime}\neq k}\left\{\boldsymbol{\mu}_{k}(\mathcal{D})-\boldsymbol{\mu}_{k}(\mathcal{D}^{\prime}_{n,k^{\prime}}))^{T}\boldsymbol{\Gamma}_{k}(\boldsymbol{\mu}_{k}(\mathcal{D})-\boldsymbol{\mu}_{k}(\mathcal{D}^{\prime}_{n,k^{\prime}})\right\},$ (10)
$\epsilon_{0}\geq\left|\ln\frac{f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}))}{f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}^{\prime}_{n,k^{\prime}}))}\right|,\forall\widetilde{\boldsymbol{\pi}}\in\mathcal{S},\forall n,k^{\prime},$ (11)
where $N_{k}$ is the number of data points in class $k$, $\epsilon_{0}>0$ and $\epsilon_{k}>0,\,1\leq k\leq K,$ are auxiliary variables.
Proof:
See Appendix A. ∎
Intuitively, Theorem 1 indicates that the global privacy guarantee (in terms of $\epsilon$) can be controlled by combining the privacy costs of each parameter release. This result gives us an explicit way to allocate the privacy budget $\epsilon_{0},\epsilon_{k},\forall k$ across the different noise additions, and then optimize the noise statistics $\{\boldsymbol{\Gamma}_{k},\gamma_{k}\}_{\forall k}$ and $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$ accordingly.
Remark 3.
Theorem 1 relies on basic DP compositions, which enable a simple closed-form expression for the KL divergence and a low-complexity solution. The choices of $\epsilon_{k}$} and $\epsilon_{0}$ provide intuitive budget trade-offs across parameters. More advanced accountants, e.g., Rényi DP [14], can tighten privacy bounds but lead to a different formulation and algorithmic structure. A full RDP-based development is an interesting direction for future work.
III-B DP-Constrained KL-divergence Minimization
We now formulate an optimization problem for the DP mechanism design. Our goal is to allocate the privacy budget $\epsilon_{0}$ and $\epsilon_{k},\forall k$, and subsequently choose the noise statistics $\{\boldsymbol{\Gamma}_{k},\gamma_{k}\}_{\forall k}$ and $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$ so as to minimize the expected KL divergence in (8), subject to a given $(\epsilon,\delta)$-DP constraint. The following theorem provides a tractable representation of the objective function in terms of these designing variables.
Proposition 1.
The expected KL divergence in (8) can be rewritten as (12), shown at the top of the next page. Here, $\psi_{d}(\cdot)$ is the multivariant digamma function. The last term depends solely on the data dimension $d$.
Proof:
See Appendix B. ∎
[…]
We propose to solve the following KL-divergence minimization problem:
[…]
$\penalty 10000\ \penalty 10000\ \text{s.t. }\max_{k}\left\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\right\}+\epsilon_{0}\leq\epsilon,$ (13b)
[…]
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \boldsymbol{\Gamma}_{k}\succeq{\bf 0},\gamma_{k}>0,\epsilon_{k}>0,\forall k,$ (13d)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \epsilon_{0}>0,{\bf F}{\bf 1}={\bf 1},[{\bf F}]_{i,j}\geq 0,\forall 1\leq i,j\leq|\mathcal{S}|.$ (13e)
Constraints (13b) and (13c) ensure the $(\epsilon,\delta)$-DP requirement, while (13d) and (13e) enforce feasibility of the precision matrices and the transition PMF. Note that (P1) is a non-convex problem and involves $\mathcal{O}(|\mathcal{S}|^{2})$ constraints, posing a challenge for large $|\mathcal{S}|$.
In (P1), we optimize the privacy budgets $\{\epsilon_{k}\}_{k=1}^{K}$ and $\epsilon_{0}$ together with the mechanism parameters. The feasibility constraints explicitly depend on problem hyperparameters such as $K$, $d$, and class sizes $\{N_{k}\}$. Consequently, smaller $N_{k}$ tightens the DP constraints and the optimizer responds by adjusting $\epsilon_{k}$ and increasing the noise power in $\boldsymbol{\Gamma}_{k}$ and $\gamma_{k}$; larger $N_{k}$ has the opposite effect. This joint treatment induces an automatic, data-aware allocation across classes and parameters, avoiding manual tuning and typically yielding a tighter KL divergence utility. We refer the reader to the numerical results in Section V for a detailed investigation of the impact of key system hyperparameters.
IV DP Mechanism Optimization
In this section, we present several modifications to (P1) that simplify the DP mechanism optimization. We then propose a low-complexity algorithm to compute a sub-optimal solution based on alternating optimization.
IV-A Problem Simplification
Recall that $|\mathcal{S}|={{N-1}\choose{K-1}}$. The dimension of $\bf F$} grows exponentially in $N$, making (P1) computationally prohibitive for large $N$. However, we note that while $\mathcal{S}$ may be large, the adjacent datasets $\{\mathcal{D}^{\prime}_{n,k^{\prime}}\}$ of $\mathcal{D}$ correspond to only a small number of possible frequency vectors $\{\boldsymbol{\pi}(\mathcal{D}^{\prime}_{n,k^{\prime}})\}$. Specifically, altering the label of exactly one data point changes exactly two entries in $\boldsymbol{\pi}(\mathcal{D})$. Motivated by this, we define the possible frequency vectors generated by all the adjacent datasets of $\mathcal{D}$ as
$\{\boldsymbol{\pi}(\mathcal{D}^{\prime}_{n,k^{\prime}})\}=\mathcal{S}^{\prime}(\mathcal{D})$
$=$ $\big\{\boldsymbol{\pi}(\mathcal{D}^{\prime})\!\in\!\mathcal{S}\!:\!\lVert\boldsymbol{\pi}(\mathcal{D}^{\prime})-\boldsymbol{\pi}(\mathcal{D})\rVert_{0}=2,\!\lVert\boldsymbol{\pi}(\mathcal{D}^{\prime})-\boldsymbol{\pi}(\mathcal{D})\rVert_{1}=\frac{2}{N}\big\}.$ (14)
It follows that $\mathcal{S}^{\prime}(\mathcal{D})\subset\mathcal{S}$ and $|\mathcal{S}^{\prime}(\mathcal{D})|=K(K-1)$. Since $|\mathcal{S}^{\prime}(\mathcal{D})|\ll|\mathcal{S}|$, to simply the problem in (P1) we parametrize the mapping PMF only over $\boldsymbol{\pi}(\mathcal{D})$ and its neighbors $\mathcal{S}^{\prime}(\mathcal{D})$. However, directly zeroing the probability outside $\{\boldsymbol{\pi}(\mathcal{D})\}\cup\mathcal{S}^{\prime}(\mathcal{D})$ yields dataset-dependent support and hence violates DP, since adjacent inputs would induce different supports. We therefore apply the following smoothing:
$f(\widetilde{\boldsymbol{\pi}}\mid\boldsymbol{\pi})=(1-\lambda)f^{\prime}(\widetilde{\boldsymbol{\pi}}\mid\boldsymbol{\pi})+\lambda/|\mathcal{S}|,\qquad\forall\widetilde{\boldsymbol{\pi}}\in\mathcal{S},$ (15)
with $\lambda\in(0,1)$ a predefined small constant, $|\mathcal{S}|={{N-1}\choose{K-1}}$, and $f^{\prime}$ supported on $\mathcal{S}^{\prime}(\mathcal{D})\cup\{\boldsymbol{\pi}\}$. The design of $f^{\prime}$ enforces nonzero mapping probability only when $\widetilde{\boldsymbol{\pi}}\in\mathcal{S}^{\prime}(\mathcal{D})$ or $\widetilde{\boldsymbol{\pi}}=\boldsymbol{\pi}(\mathcal{D})$, while $f$ attains dataset-independent support $\mathcal{S}$. By post-processing, if $f^{\prime}$ is $(\epsilon_{0},0)$-DP then $f$ is also $(\epsilon_{0},0)$-DP. Hence, the corresponding transition matrix $\bf F$} has non-zero entries only in a submatrix ${\bf F}^{\prime}\in[0,1]^{(|\mathcal{S}^{\prime}(\mathcal{D})|+1)\times(|\mathcal{S}^{\prime}(\mathcal{D})|+1)}$, whose rows and columns are indexed by $\boldsymbol{\pi}(\mathcal{D})$ and the elements of $\mathcal{S}^{\prime}(\mathcal{D})$. This strategy reduces the number of free variables in $\bf F$} from $\mathcal{O}(e^{N})$ to $\mathcal{O}(K^{4})$.
Under this restriction, the original constraint (11) becomes:
$\epsilon_{0}\geq\left|\ln\frac{f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}))}{f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}^{\prime}))}\right|,\forall\widetilde{\boldsymbol{\pi}}\in\mathcal{S}^{\prime}(\mathcal{D})\cup\{\boldsymbol{\pi}(\mathcal{D})\},\forall\boldsymbol{\pi}^{\prime}\in\mathcal{S}^{\prime}(\mathcal{D}).$ (16)
Fixing an indexing order for $\{\boldsymbol{\pi}(\mathcal{D})\}\cup\mathcal{S}^{\prime}(\mathcal{D})$, let $\boldsymbol{\pi}(\mathcal{D})$ correspond to the $j^{\star}$-th column of $\bf F$$^{\prime}$. Then (16) is equivalent to
$e^{-\epsilon_{0}}F_{i,j}^{\prime}\leq F_{i,j^{\star}}\leq e^{\epsilon_{0}}F_{i,j}^{\prime},\forall j\neq j^{\star},\forall 1\leq i\leq|\mathcal{S}^{\prime}(\mathcal{D})|+1.$ (17)
Next, we simplify the constraints involving $\boldsymbol{\Gamma}_{k}$ by switching to its inverse $\boldsymbol{\Gamma}_{k}^{-1}$. Using the Schur complement on $\boldsymbol{\Gamma}_{k}^{-1}$, we have
$\boldsymbol{\Gamma}_{k}\succeq{\bf 0}\Leftrightarrow\boldsymbol{\Gamma}_{k}^{-1}\succeq{\bf 0},$ (18)
[…]
Therefore, we can solve for $\boldsymbol{\Gamma}_{k}^{-1}$ directly, ensuring all relevant constraints are convex in $\boldsymbol{\Gamma}_{k}^{-1}$.
Gathering all the above modifications, we obtain a simplified optimization problem:
$(\text{P2}):$ $\min_{\begin{subarray}{c}\{\gamma_{k},\boldsymbol{\Gamma}_{k}^{-1},\epsilon_{k}\}_{k=1}^{K}\\ {\bf F}^{\prime},\epsilon_{0}\end{subarray}}\sum_{\widetilde{\boldsymbol{\pi}}}f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}))g(\widetilde{\boldsymbol{\pi}},\{\gamma_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{\forall k})$ (20a)
$\penalty 10000\ \penalty 10000\ \text{s.t. }\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}+\epsilon_{0}\leq\epsilon,\forall k$ (20b)
[…]
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \boldsymbol{\Gamma}_{k}^{-1}\succeq{\bf 0},\gamma_{k}>0,\epsilon_{k}>0,\forall k,$ (20d)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \epsilon_{0}>0,{\bf F}^{\prime}{\bf 1}={\bf 1},[{\bf F}^{\prime}]_{i,j}\geq 0,\forall i,j,$ (20e)
where $g(\widetilde{\boldsymbol{\pi}},\{\gamma_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{\forall k})$ is defined in (12).
Because the feasible region of (P2) is a subset of that for (P1), the solution to (P2) is generally suboptimal but always feasible for (P1). Nonetheless, (P2) reduces the complexity substantially, down to $\mathcal{O}(K^{4}+NK^{2})$ convex constraints compared to the exponential size of $\bf F$} in (P1).
IV-B Proposed Solution to (P2)
Although Problem (P2) is now significantly simplified to a reduced number of convex constrains, the non-convex nature of the objective still makes its optimal solution intractable. In this section, we propose a sub-optimal solution that optimizes the variables in an alternating fashion. The proposed approach first initializes a feasible guess for $\epsilon_{0}$ and $\epsilon_{k},\forall k$, and then alternately solves for $\{\gamma_{k}\}_{k=1}^{K}$, $\{\epsilon_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{k=1}^{K}$, and $\{{\bf F}^{\prime},\epsilon_{0}\}$ until convergence. The details are listed as follows.
  1. 1.
    Update $\{\gamma_{k}\}_{k=1}^{K}$: Fixing the values of the other variables, the optimization of $\{\gamma_{k}\}_{k=1}^{K}$ can be decomposed into $K$ independent one-dimensional optimization subproblems as
    $\min_{\gamma_{k}>0}\penalty 10000\ d\gamma_{k}+\frac{(d+1)\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1})}{\gamma_{k}}$
    $\text{ s.t. }\penalty 10000\ \gamma_{k}\leq\frac{2N_{k}(\epsilon-\epsilon_{0}-\epsilon_{k})}{3}.$ (21)
    By using the first-order optimality condition, it can be verified that the optimal solution to $\gamma_{k}$ is given by
    $\gamma_{k}=\min\left\{\frac{d+1}{d}\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1}),\frac{2N_{k}(\epsilon-\epsilon_{0}-\epsilon_{k})}{3}\right\},\forall k.$ (22)
  2. 2.
    Update $\{\epsilon_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{k=1}^{K}$: Fixing the other varaibles, the optimziation of $\{\epsilon_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{k=1}^{K}$ can be decomposted into the following $K$ independent convex subproblem:
    $\min_{\boldsymbol{\Gamma}_{k}^{-1},\epsilon_{k}^{2}}\penalty 10000\ \penalty 10000\ \mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1}\boldsymbol{\Gamma}_{k}^{-1})$
    $\text{s.t. }\boldsymbol{\Gamma}_{k}^{-1}\succeq{\bf 0},0<\epsilon_{k}^{2}\leq(\epsilon-\epsilon_{0}-\frac{3\gamma_{k}}{2N_{k}})^{2},$
    $\begin{bmatrix}\epsilon_{k}^{2}/(2\log(2/\delta))&\boldsymbol{\mu}_{k}(\mathcal{D})-\boldsymbol{\mu}_{k}(\mathcal{D}^{\prime}_{n,k^{\prime}})\\ \boldsymbol{\mu}_{k}^{T}(\mathcal{D})-\boldsymbol{\mu}_{k}^{T}(\mathcal{D}^{\prime}_{n,k^{\prime}})&\boldsymbol{\Gamma}_{k}^{-1}\end{bmatrix}\succeq{\bf 0},\forall n,k^{\prime}.$ (23)
    This is a semi-definite program (SDP) and can be solved by off-the-shelf solvers such as CVX [15]. The precision matrix $\boldsymbol{\Gamma}_{k}$ can be obtained by the inverse of $\boldsymbol{\Gamma}_{k}^{-1}$ after the algorithm converges.
  3. 3.
    Update $\bf F$$^{\prime}$ and $\epsilon_{0}$: Fixing $\{\gamma_{k},\epsilon_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{k=1}^{K}$, the optimization of $\bf F$$^{\prime}$ and $\epsilon_{0}$ can be recast as
    $\min_{{\bf F}^{\prime},\epsilon_{0}}\penalty 10000\ \sum_{i=1}^{|\mathcal{S}^{\prime}(\mathcal{D})|+1}g_{i}F^{\prime}_{i,j^{\star}}$
    $\penalty 10000\ \penalty 10000\ \text{s.t. }\epsilon_{0}\leq\epsilon-\max_{k}\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\}$
    $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ e^{-\epsilon_{0}}F_{i,j}^{\prime}\leq F_{i,j^{\star}}\leq e^{\epsilon_{0}}F_{i,j}^{\prime},\forall j\neq j^{\star},\forall i.$
    $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \epsilon_{0}>0,{\bf F}^{\prime}{\bf 1}={\bf 1},[{\bf F}^{\prime}]_{i,j}\geq 0,\forall i,j.$ (24)
    Here, $g_{i}=g(\widetilde{\boldsymbol{\pi}},\{\gamma_{k},\boldsymbol{\Gamma}_{k}^{-1}\}_{\forall k})$ for $\widetilde{\boldsymbol{\pi}}$ that corresponds to the $i$-th row of $\bf F$$^{\prime}$ and $j^{\star}$ is the index where $\boldsymbol{\pi}(\mathcal{D})$ correspond to the $j^{\star}$-th column of $\bf F$$^{\prime}$.
    The solution of (24) is given by the following proposition.
    Proposition 2.
    The optimal solution to (24) is given by
    $\epsilon_{0}$ $=\epsilon-\max_{k}\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\},$ (25)
[…]
    $F_{i,j}^{\prime}$ $=\left\{\begin{aligned} e^{\epsilon_{0}}F_{i,j^{\star}}^{\prime},&\text{ if }g_{i}\geq 0\\ e^{-\epsilon_{0}}F_{i,j^{\star}}^{\prime},&\text{ otherwise}\\ \end{aligned}\right.,\forall j\neq j^{\star}.$ (27)
    Proof:
    See Appendix C. ∎
    1:Input: $\epsilon$, $\delta$, $N$, $d$, $\boldsymbol{\pi}(\mathcal{D})$, $\{\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}$, $\{\boldsymbol{\mu}_{k}(\mathcal{D}_{n,k^{\prime}})\}$, $\mathcal{S}^{\prime}(\mathcal{D})$.
    2: Initialize $\epsilon_{0}=\epsilon_{k}=\epsilon/3,\forall k$.
    3:for $\mbox{iter}=1,2,\cdots,I$ do
    4:for $k=1,2,\cdots,K$ in parallel do
    5:  Update $\gamma_{k}$ by (22);
    6:  Update $\boldsymbol{\Gamma}_{k}^{-1}$ and $\epsilon_{k}$ by solving (23);
    7:end for
    8:  Update $\bf F$} and $\epsilon_{0}$ by (25)–(27);
    9:if $|Obj(\cdot;\mbox{iter})-Obj(\cdot;\mbox{iter}-1)|\leq 10^{-3}$, early stop;
    10:end for
    11: Compute $\boldsymbol{\Gamma}_{k}$ by the inverse of $\boldsymbol{\Gamma}_{k}^{-1}$$\forall k$;
    12: Draw $\{{\bf w}_{k},{\bf W}_{K}\}$ and $\widetilde{\boldsymbol{\pi}}$ by (5)–(6) and $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$.
    13:Output: $\widetilde{\boldsymbol{\pi}}$ and $\{\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{\forall k}$.
    Algorithm 1 The proposed DP mechanism.
    The overall procedure is summarized in Algorithm 1, shown at the top of this page. In Step 9, we include an early-stopping criterion when the change in the objective value of (20a) between two consecutive iterations falls below a small threshold $10^{-3}$. In practice, convergence is typically achieved within $10$ iterations. Since the objective is bounded below by zero and each iteration produces a non-increasing sequence of objectives, convergence of Algorithm 1 is guaranteed.
    Remark 4 (Why the proposed mechanism set matter?).
    The colored Gaussian, Wishart, and randomized-mapping mechanisms admit clean DP bounds that compose directly, while their joint effect yields a simplified KL divergence expression. This structure is essential for tractable updates: the mean-noise covariances $\{\boldsymbol{\Gamma}_{k}\}$ and Wishart scales $\{\gamma_{k}\}$ enter convex subproblems, and the mapping probabilities $\bf F$$^{\prime}$ satisfy linear constraints. The alternating scheme thus provides (i) fast convergence and (ii) interpretable privacy–utility control via dynamic reallocation of $\{\epsilon_{k}\}$ and $\epsilon_{0}$ across parameters.
    Remark 5 (Complexity Analysis).
    The primary computational cost of Algorithm 1 arises from solving the SDP problem in (23) and computing $\bf F$} via (26) and (27). According to [16], the SDP in (23) involves a $d\times d$ matrix variable and $\mathcal{O}(NK^{2})$ constraints. Solving the problem via an interior-point method incurs a complexity of $\mathcal{O}(N^{3.5}K^{6})$ in the regime of $N\gg d$ [16]. Meanwhile, the computation in (26) and (27) requires $\mathcal{O}(K^{4})$ floating-point operations. Consequently, the overall complexity of Algorithm 1 is on the order of $\mathcal{O}(N^{3.5}K^{6})$. Notably, this complexity grow in a polynomial order with the dataset size $N$, due to our restriction that the probability mapping matrix $\bf F$} only takes non-trivial entries for $\boldsymbol{\pi}(\mathcal{D})$ and its adjacent frequency vectors, thus avoiding an exponential increase with $N$.
    IV-C Adaptions to Other DP Definitions
    As stated in Definition 1, we consider label-level DP, where adjacent datasets differ by flipping exactly one label. For a dataset with $N$ points, all label-level adjacent datasets can be enumerated explicitly, which leads to the data-dependent bounds in Theorem 1.
    On the other hand, we show here that this label-level DP setting is technically more challenging than record-level DP, and our analysis and method in Algorithm 1 adapt readily to record-level DP under a uniform feature-space bound. Consider DP defined with an adjacent dataset that adds, removes, or replaces one data point, and assume a uniform bound on features,
    $\lVert{\bf x}_{n}\rVert_{2}\leq B,\ \forall n,$ (28)
    for some constant $B<\infty$.
    We discuss the adaption to three sub-cases for record-level DP:
    • $\bullet$
      Removing one data point. All arguments from the label-DP case carry over, except for the adjacent mixture-weight set. Under label flips, one sample moves between two classes, changing two coordinates of $\boldsymbol{\pi}$. The set of adjacent weights $\widetilde{\boldsymbol{\pi}}$ for any $\mathcal{D}^{\prime}$ is given by (14) with K(K1)K(K-1) elements.
      In contrast, under remove-one record-level DP, only one coordinate of $\boldsymbol{\pi}$ changes. The set of adjacent mixture weights in this case is given by
      $\{\boldsymbol{\pi}(\mathcal{D}^{\prime}_{n,k^{\prime}})\}=\mathcal{S}^{\prime}(\mathcal{D})$
      $=$ $\big\{\boldsymbol{\pi}(\mathcal{D}^{\prime})\!\in\!\mathcal{S}\!:\!\lVert\boldsymbol{\pi}(\mathcal{D}^{\prime})-\boldsymbol{\pi}(\mathcal{D})\rVert_{0}=1,\!$
      $\lVert(N-1)\boldsymbol{\pi}(\mathcal{D}^{\prime})-N\boldsymbol{\pi}(\mathcal{D})\rVert_{1}=1\big\},$ (29)
      which has only $K$ elements. With (29), all derivations and designs remain intact, except that the constraint in (17) is now evaluated over the $K$ elements in (29). This case thus induces fewer constraints.
    • $\bullet$
      Changing one feature value. Here an adjacent dataset alters one entry of some ${\bf x}_{n}$ while keeping labels fixed, so all adjacent datasets share the same mixture weight $\boldsymbol{\pi}$. Therefore the release of $\boldsymbol{\pi}$ is inherently DP and does not require randomization. We have $\widetilde{\boldsymbol{\pi}}=\boldsymbol{\pi}$, and we may drop the variables $\bf F$} and $\epsilon_{0}$ from the optimization in Problem P11.
      With the feature bound in (28), we also obtain a similar closed-form DP bound for the Gaussian mechanism for releasing class sample means. For an adjacent dataset $\mathcal{D}^{\prime}$ that changes one feature value,
      $\lVert\boldsymbol{\mu}_{k}(\mathcal{D})-\boldsymbol{\mu}_{k}(\mathcal{D}^{\prime})\rVert\leq\frac{2B}{N_{k}},$ (30)
      where $N_{k}$ is the number of points in class $k$. Therefore, $\widetilde{\boldsymbol{\mu}}_{k}$ attains $(\epsilon_{k},\delta)$-DP provided (cf. [4, Eq. (14)])
      $\frac{\epsilon_{k}^{2}N_{k}^{2}}{2\ln(2/\delta)}\geq 4B^{2}\lVert\boldsymbol{\Gamma}_{k}\rVert_{2}.$ (31)
      where $\lVert\boldsymbol{\Gamma}_{k}\rVert_{2}$ is the spectral norm. This replaces the data-dependent bound in (31) with a (potentially looser) data-independent one. Consequently, we obtain $K$ constraints instead of the N(K1)N(K-1) constraints in (19):
      $\boldsymbol{\Gamma}_{k}^{-1}-\frac{8B\ln(2/\delta)}{N_{k}^{2}\epsilon_{k}^{2}}{\bf I}_{d}\succeq{\bf 0},\ \forall k.$ (32)
      Our algorithm remains applicable; only the update of $\boldsymbol{\Gamma}_{k}^{-1}$ in (23) is replaced by a simpler SDP problem with fewer constraints. This variant is substantially simpler than the setting for the label-level DP since $\boldsymbol{\pi}$ needs no additional protection.
    • $\bullet$
      Adding one bounded data point. The adjacent mixture-weight set is adjusted similar to (29), again yielding $K$ constraints in place of the K(K1)K(K-1) constraints in (17). Moreover, inserting a point into one class (say, Class $k_{0}$) introduces one additional constraint of the form (32) for that class, i.e., one extra constraint in (23).
    Taken together, these cases show that the label-level DP defined in Definition 1 is the most technically demanding. Once this case is handled, adapting the approach to record-level notions entails only minor, principled modifications. Accordingly, we retain label-level DP as the default setup.
    Complexity and a scalable variant. The SDP in (23) is constrained by the $\mathcal{O}(N)$ data-dependent DP conditions in (19), which leads to a worst-case complexity growing with $N$. For large datasets, we offer a scalable alternative based on a uniform feature bound in (28), which yields the per-class DP condition in (32). This replacement reduces the number of constraints from N(K1)N(K-1) to $K$, leading to an overall complexity $\mathcal{O}(K^{6})$ that is independent of $N$. While this bound is more conservative than the data-dependent one, our experiments indicate comparable qualitative privacy–utility trends.
    V Experimental Results
    In this section, we present the experimental results to evaluate the performance of the proposed approach.
    V-A Results on Synthetic Data
    We begin by assessing the performance of the proposed method on a synthetic dataset drawn from a GMM. Unless otherwise specified, the simulation setup is as follows. First, we generate a ground-truth GMM with the following parameters: the class frequency vector is drawn from a Dirichlet distribution with $K=5$ categories and concentration parameters equal to one; the mean of each component is drawn from a uniform distribution within the range $10,10]^{d}$, with data dimension $d=3$; and the covariance of each component is drawn from a Wishart distribution with $d+1$ degrees of freedom and a scale matrix ${\bf I}_{d}$.
    Next, we use the GMM with these parameters to generate $N=1000$ independent and identically distributed (i.i.d.) samples to form the input dataset $\mathcal{D}$. This dataset is then fitted to an empirical GMM $p({\bf x},y)$ as defined in (1) using the approach outlined in (2). We then apply the proposed DP mechanism to compute a differentially private GMM $\widetilde{p}({\bf x},y)$ using Algorithm 1, with a predefined $(\epsilon,\delta)$-DP requirement. Unless otherwise specified, we set $\delta=10^{-5}$ and adjust $\epsilon$ to control the privacy level. The accuracy of the released model is measured by the KL divergence between the released GMM and the non-private model, i.e., $KL\left(\widetilde{p}({\bf x},y)|p({\bf x},y)\right)$.
    We compare the performance of the proposed method against the following existing DP mechanisms:
    • $\bullet$
      i.i.d. Laplace mechanism [6, Algorithm 2]: This baseline method adds i.i.d. Laplace noise to the estimated parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ and then projects the perturbed parameters back onto the feasible set. As demonstrated in [6], the noise added is determined by the $\ell_{1}$ sensitivity of the individual parameters. This mechanism guarantees $(\epsilon,0)$-DP, making it a stronger privacy condition and sufficient for achieving $(\epsilon,\delta)$-DP for any $\delta>0$.
    • $\bullet$
      i.i.d. Gaussian mechanism [2]: In this baseline method, i.i.d. Gaussian noise is added to the estimated parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$, followed by a projection of the perturbed parameters back onto the feasible set to ensure the release satisfies the $(\epsilon,\delta)$-DP constraint. The variance of the added noise is determined by the $\ell_{2}$ sensitivity of the parameters. Based on the norm inequality, the $\ell_{2}$ sensitivity is bounded by the corresponding $\ell_{1}$ sensitivity, which is analytically expressed in [6].
    • $\bullet$
      Colored Gaussian mechanism [4, Eq. (12)]: Originally designed to enhance DP for $K$-Means clustering, this method is adapted here to protect the mean vectors of the GMM components. Unlike the standard i.i.d. Gaussian mechanism, colored Gaussian noise is added to the mean components $\{{\bf w}_{k}\}_{k=1}^{K}$ according to the approach in (5). The covariance matrix $\boldsymbol{\Gamma}_{k}^{-1},\forall k,$ is optimized by minimizing the MSE between the perturbed and true means, subject to the $(\epsilon,\delta)$-DP constraint.
    • $\bullet$
      Rank-1 Singular Multivariate Gaussian (R1SMG) [17]: This baseline calculates a Gaussian noise according to [17, Eq. (9)] and adds the noise to the class sample means $\{\boldsymbol{\mu}_{k}\}$. The covariance and weight mechanisms remain as in the above baselines.
    For a fair comparison, we ensure that all the above DP mechanisms all satisfy the $(\epsilon,\delta)$-DP with the same $\epsilon$ and $\delta$ under the same accounting framework. Note that Laplace mechanisms typically achieve $(\epsilon,0)$-DP, which is a stronger condition and, therefore, sufficient to guarantee $(\epsilon,\delta)$-DP for any $\delta>0$. All results are averaged over $100$ Monte Carlo trials unless otherwise specified. For our approach, we set the parameters $\lambda=10^{-3}$.
    Figure 1: Average KL divergence and confidence interval for the proposed method across $200$ Monte Carlo trials. The privacy level is represented by the value of $\epsilon$.
    Figure 2: KL divergence versus the privacy level in terms of the value of $\epsilon$.
    We first plot the performance of the proposed method with $200$ Monte Carlo trials. Figure 1 shows the mean value of the KL divergence as well as the 95% confidence interval (error bar) under different privacy levels. The simulation is run in MATLAB R2025b on a MacOS laptop (Apple M4, 10-core CPU, 32 GB memory). We set the random seed to $42$, the smoothing parameter to $\lambda=10^{-3}$, and the early-stopping threshold to $10^{-3}$. SDP subproblems are solved in CVX [15] using MOSEK as the default solver with default solver hyperparameters. In practice, the proposed algorithm typically terminates within $10$ iterations and each trial finishes in $10$ seconds.
    We examine the utility-privacy trade-off for releasing GMMs. Figure 2 plots the average KL divergence of the DP method under varying privacy constraints for different algorithms, as the value of $\epsilon$ changes. As expected, a larger $\epsilon$ results in more artificial noise, leading to greater distortion in the GMM parameters, which is reflected in a higher KL divergence. The proposed approach consistently achieves a significantly smaller KL divergence than all the other baselines, thanks to its capability to directly minimize this utility measure. The performance of R1SMG is consistent with the calibration in [16] for small ambient (or retained) dimension $d$, where the noise variance factor in the R1SMG algorithm (cf. [16, Eq. (9)]) becomes large. In the following simulations, we drop this baseline and compare our method with other mechanisms.
    Figure 3: KL divergence versus the data size $N$ with $\epsilon=1$.
    Figure 4: KL divergence versus the number of classes $K$, where the total data size is fixed to $N=1000$.
    Next, we fix $\epsilon=1$ and examine the KL divergence performance as a function of the total number of data points $N$, as shown in Figure 3. We see that the KL divergence decreases as $N$ increases under the same DP constraint. Intuitively, DP assesses the probability of inferring the presence or properties of a particular data point in the entire dataset. Since the data points are i.i.d. and drawn from the same GMM distribution, a larger dataset naturally has a greater capacity to obscure individual data points. As a result, it requires less noise to achieve the same level of privacy. This observation aligns with the DP analysis in Section III.
    Another important hyperparameter that affects the performance of differentially private GMMs is the number of data classes $K$. By fixing $N=1000$ and $\epsilon=1$, we plot the KL divergence performance against varying values of $K$ in Figure 4. As shown in Theorem 1, the privacy loss from releasing the component mean ${\bf w}_{k}$ and covariance ${\bf W}_{k}$ for each class $k$ is independently determined by the statistics of the added noise. The overall privacy loss for releasing $\{\widetilde{\bf w}_{k},\widetilde{\bf W}_{k}\}_{\forall k}$ is governed by the maximum privacy loss for any class $k$, making it a non-decreasing function of $K$. Specifically, the relationship between the KL divergence and $K$ is governed by the following two factors:
    • $\bullet$
      Additive structure of the KL expression. The closed-form KL divergence we use in (12) is additive across components. Holding the per-component terms comparable, adding more components increases the total sum, so the KL divergence tends to be non-decreasing in $K$. Intuitively, a finer mixture with more modes (larger $K$) produces more heterogeneous fitted distributions, which enlarges the distance between the private and non-private GMMs.
    • $\bullet$
      Stronger DP perturbation. With $K$ larger and NN fixed, the smallest class size $N_{k}$ decreases. The privacy losses for releasing $\{\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}$ therefore worsen: the Wishart privacy loss bound scales as $\tfrac{3\gamma_{k}}{2N_{k}}$ (cf. (9)), and the Gaussian privacy loss for sample means also increases as $N_{k}$ shrinks (cf. (10)).
      Furthermore, with bounded feature vectors as shown in (28), the privacy loss of the mean-release in (31) shows that, to maintain the same $\epsilon_{k}$ with smaller $N_{k}$, one must increase the noise power in $\boldsymbol{\Gamma}_{k}$ and $\gamma_{k}$. Since our KL expression grows with these noise levels, the resulting KL increases with $K$.
    This analysis is confirmed by the results in Figure 4. However, the increase in KL divergence for the proposed approach is much less sensitive to $K$ than in the baseline methods, highlighting the robustness of the proposed approach.
    Finally, we evaluate the KL divergence with respect to the data dimension $d$ in Figure 5. From Theorem 1 and Proposition 1, it can be seen that under the same statistics, both the KL divergence and the privacy loss increase with $d$. Intuitively, due to the curse of dimensionality, data points become sparsely distributed in high-dimensional spaces, increasing their susceptibility to privacy risks as adjacent datasets become easier to distinguish. As a result, more perturbations are required to maintain the same level of DP. As shown in Figure 5, the accuracy of all DP mechanisms significantly deteriorates as $d$ increases, demonstrating the sensitivity of differentially private GMMs to data dimensionality.
    Figure 5: KL divergence versus the dimension of the data points $d$ with $K=5$ and $N=1000$.
    V-B Results on the UCI Machine Learning Dataset
    Figure 6: KL divergence versus the value of $\epsilon$ for releasing the GMM fitting to the Iris dataset.
    Figure 7: Marginal probability density of GMMs and empirical histograms of the true values fitted to the Iris dataset for the first feature, i.e., the first entry of the data points $\{{\bf x}_{n}\}_{n=1}^{N}$. Each subplot displays the conditional density for one class.
    In this section, we evaluate the performance of the proposed DP method using real-world multi-class datasets. We employ the Iris dataset from the UCI Machine Learning Repository [18], which is a widely recognized benchmark for classification tasks. The dataset consists of $K=3$ classes, each representing a type of iris plant, with a total of $N=150$ instances. Each instance is characterized by a feature vector of $d=4$ dimensions, measuring the sepal length, sepal width, petal length, and petal width. These features are normalized to have zero mean and unit variance. Notably, the first class (Iris-Setosa) is linearly separable from the other two, while the latter two classes (Iris-Versicolor and Iris-Virginica) exhibit overlapping distributions, making them harder to distinguish.
    We fit the dataset to an empirical GMM $p({\bf x},y)$, as defined in (1), using the method outlined in (2). Subsequently, we apply the proposed DP mechanism to compute a differentially private GMM $\widetilde{p}({\bf x},y)$, with a predefined $(\epsilon,\delta)$-DP constraint, where we choose $\epsilon=2$ and $\delta=10^{-5}$. Figure 7 visualizes the marginal probability density for the first feature (normalized sepal length) for each class. The gray bars represent the empirical histograms of the true feature values, and the black dashed lines represent the fitted GMM density in (2) without applying DP. The results show that the GMM produced by our method aligns more closely with the true values compared to the baseline method.
    Moreover, Figure 6 illustrates the KL divergence between the differentially private GMM and the non-private model, under varying levels of privacy determined by $\epsilon$. Our approach exhibits a superior utility-privacy trade-off, achieving a smaller KL divergence under the same DP constraints. The results in Figures 6 and 7 further justify the use of KL divergence as a utility metric, since a smaller divergence indicates a better fit to the ground truth, demonstrating the effectiveness of our method in preserving data accuracy while maintaining privacy.
    V-C Results on Electricity Load Profile Data
    Data class # of data points
    Class 11 $687$
    Class 22 $428$
    Class 33 $102$
    Class 44 $147$
    Class 55 $14$
    Class 66 $31$
    Total $1409$
[…]
    Table I: The number of data points in each class of the AMI data.
    Figure 8: KL divergence versus the value of $\epsilon$ for releasing the GMM fitting to the AMI load profile data.
    Figure 9: Entry-wise marginal probability density of GMMs fitted to the AMI load profile data for each class. The subplots in each column correspond to the marginal density of one class. In these plots, the gray bars represent the empirical histograms of the true values, the blue curves represent the marginal density of the GMM computed using the proposed DP method, and the red curves represent the marginal density of the non-private GMMs computed using (2).
    We explore the application of differentially private GMMs in real-world scenarios, specifically within the context of power systems. Electricity load demand profiles, which represent the amount of electricity consumption over time, are essential for system analysis, anomaly detection, and fault diagnosis in power systems. These load profile data are typically collected by a utility provider and shared with an analyst for further analysis. It has been shown in [4] that the logarithmic values of load demand profiles fit well to GMMs. However, recent studies have highlighted significant privacy risks associated with releasing load profile data or their statistics, motivating the release of differentially private GMMs for load profile synthesis and analysis.
    Here, we examine the performance of fitting and releasing GMMs on real-world load profile data under DP constraints. Following [4], we utilize a dataset from a real-world AMI system, which includes data from $1409$ houses spread across 12 distribution circuits in California, USA, yielding $N=1409$ samples. Each sample represents the half-day electricity consumption profile of a single house, with measurements taken hourly over $d=12$ consecutive hours. These data points are then transformed by taking the logarithm of the power consumption. We then apply $K$-Means clustering [19] to group the data into $K=6$ classes, representing different types of electricity consumers (e.g., residential, commercial, agricultural). Table I summarizes the number of data points in each class.
    We proceed by fitting the dataset to a GMM using the method in (2) and adding artificial noise to achieve $(\epsilon,\delta)$-DP. Figure 8 shows the KL divergence between the released GMM and the non-private model, plotted against different values of $\epsilon$, with $\delta=10^{-5}$ fixed. The results reveal that our proposed method significantly improves the utility in terms of KL divergence compared to the baseline algorithms, effectively balancing the privacy-utility trade-off.
    Next, we fix $\epsilon=2$ and plot the marginal density of the released GMM for each dimension and class in Figure 9. For Classes 1-4, the density values closely align with the fitted GMM without DP (represented by the red curves) and the ground-truth histograms (represented by the gray bars). However, for Classes 5 and 6, the release of their parameters is more sensitive due to the limited data available, requiring significantly more noise to maintain DP. As a result, the density values from the proposed approach diverge from those of the non-private models in some dimensions for these classes. This highlights the sensitivity of differentially private GMMs to sample size, particularly when fewer data points are available.
    V-D Results on Image Classification
    Figure 10: KL divergence versus the value of $\epsilon$ for releasing the GMM fitting to the MNIST training data using the GMM classifier.
    GMMs are widely adopted models for classification. In this section we explore their use in a differentially private setting for image classification. We consider the well-known MNIST dataset [20], which contains $28\times 28=784$ grayscale images of handwritten digits. We randomly sample $N=5000$ training images and reduce the dimensionality from $784$ to $d=5$ using principal component analysis [21] to mitigate the curse of dimensionality. Next, we fit these data to a $10$-class GMM using the approach in (2) and subsequently apply the DP mechanism to enforce an $(\epsilon,\delta)$-DP constraint with $\delta=10^{-5}$.
    Figure 10 plots the KL divergence between the differentially private GMM and the non-private model as a function of the privacy parameter $\epsilon$. The proposed approach achieves a significantly lower divergence, indicating that the released GMM classifier closely approximates the non-private model. Furthermore, we evaluate classification performance by predicting the labels of $10^{4}$ test images, where the class for each test vector is determined via maximum likelihood estimation by using the GMM. Fig. 11 plots the test accuracy of different classifiers, measured by the fraction of correctly predicted labels in the range of $[0,1]$. We see that our method achieves an accuracy close to the non-private baseline, while the DP baseline approaches suffer from high distortion due to added noise. These results validate the effectiveness of using KL divergence as a utility metric, as a smaller divergence corresponds to a closer match between the training and testing data distributions, thereby enhancing classification performance.
    The KL divergence grows with the ambient dimension $d$, both because the KL divergence expression aggregates across dimensions and because DP calibration induces larger noise in higher dimensions. Combining Figures 5 and 10, we observe that PCA-based dimensionality reduction preserves task-relevant variability while lowering the required noise scale, thereby mitigating the curse of dimensionality and reducing the overall noise at a fixed $(\epsilon,\delta)$-DP level.
    The above simulation demonstrates the accuracy of DP-GMM for direct classification. In the following, we extend the simulation to an end-to-end use case where a DP-GMM is queried to synthesize training data for a downstream classifier on MNIST. Concretely, the data owner samples $N=5000$ training images from MNIST and releases the $10$-class GMM parameters $\{\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k},\widetilde{\boldsymbol{\pi}}\}_{k=1}^{10}$ under an $(\epsilon,\delta)$-DP budget with $\delta=10^{-5}$ using our method. We apply the same PCA pipeline as in Section V-A (retain $d=5$ principal components) and query the DP-GMM to generate $N=5000$ synthetic samples class-conditionally so that the class histogram is preserved. By the post-processing theorem, the synthetic dataset remains $(\epsilon,\delta)$-DP with respect to the original training set.
    For the downstream model, we train a $3$-layer multi-layer perceptron (MLP), fully connected $d\to 256\to 128\to 10$ with ReLU activations and a softmax loss using stochastic gradient descent (SGD) with mini-batch size $128$, learning rate $0.01$, and a single epoch per training run. Testing is performed on the held-out ground-truth MNIST test set. Figure 12 reports test accuracy versus training iteration for synthetic datasets generated at different privacy levels. We also include two references: (i) a model trained on the non-private ground-truth training data (black curve) and (ii) a model trained on non-private GMM-synthetic data (red curve). The curves exhibit a clear privacy–utility trade-off: more stringent privacy (smaller $\epsilon$) induces larger KL divergence in the DP-GMM fit and reduces downstream accuracy. At the same time, the proximity between the black and red curves confirms that GMM-based synthetic data can train an effective classifier, and the consistent ordering across $\epsilon$ validates the suitability of KL divergence as a utility proxy for this task.
    Figure 11: Test accuracy versus the value of $\epsilon$ over the MNIST testing data using the GMM classifier.
    Figure 12: Test accuracy over the MNIST testing data using the MLP model trained by synthetic data.
    V-E Results on Record-Level DP
    As discussed in Section IV-C, the proposed method extends to record-level DP under bounded feature vectors. Here we evaluate this setting, where adjacency is defined by changing the features of a single data point. Using the synthetic dataset in Section V-A, we fit a GMM with an additional feature-clipping step: for each data point ${\bf x}_{n}$, we apply per-record clipping ${\bf x}_{n}\leftarrow{\bf x}_{n}\cdot\min\{B/\lVert{\bf x}_{n}\rVert_{2}\}$ so that (28) holds. We then compute $(\epsilon,\delta)$-DP and adapt our algorithm exactly as described in Section IV-C.
    Figure 13 reports the privacy–utility trade-off for varying clipping thresholds $B$ at $\epsilon=1$. Consistent with (32), achieving a fixed per-class privacy budget $\epsilon_{k}$ with a smaller $B$ requires larger Gaussian noise, which in turn increases the KL divergence for the same algorithmic settings. This trend matches the curves in Figure 13 and supports the effectiveness of the proposed method for record-level DP preservation.
    Figure 13: KL divergence versus the clipping threshold $B$ for record-level DP preservation.
    VI Conclusions
    In this paper, we addressed the privacy protection challenge associated with releasing the parameters of GMMs, including mixture weights, component means, and covariance matrices. We proposed the use of KL divergence as a utility metric to quantify the accuracy of the released GMM, which effectively captures the combined impact of noise perturbation on individual parameters. We then introduced a DP mechanism designed to protect the privacy of GMM parameters. Our analysis reveals the impact of privacy budget allocation and noise statistics on DP and also offers a tractable expression for evaluating the KL divergence utility. Building on this analysis, we formulated and solved an optimization problem that minimizes the KL divergence between the released and original models, subject to a given $(\epsilon,\delta)$-DP constraint. Experimental results on both synthetic and real-world datasets demonstrate the effectiveness of our approach, highlighting its superior performance in achieving a balance between privacy and accuracy.
    Appendix A Proof of Theorem 1
    We first characterize the DP of releasing $\{\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}$ using the following established results.
    Lemma 1 (Multivariant Gaussian mechanism [4]).
    Given $\mathcal{D}$, the mechanism for publishing $\widetilde{\boldsymbol{\mu}}_{k}$ in (5) satisfies $(\epsilon_{k},\delta)$-DP given that the following inequality holds for any adjacent dataset $\mathcal{D}^{\prime}_{n,k^{\prime}}$:
    $\frac{\epsilon_{k}^{2}}{2\log(2/\delta)}$
    $\geq\sup_{n,k^{\prime}}\left\{\boldsymbol{\mu}_{k}(\mathcal{D})-\boldsymbol{\mu}_{k}(\mathcal{D}^{\prime}_{n,k^{\prime}}))^{T}\boldsymbol{\Gamma}_{k}(\boldsymbol{\mu}_{k}(\mathcal{D})-\boldsymbol{\mu}_{k}(\mathcal{D}^{\prime}_{n,k^{\prime}})\right\}.$
    Proof:
    See [4, Theorem 3]. ∎
    Lemma 2 (Wishart mechanism [10]).
    The mechanism for publishing $\widetilde{\boldsymbol{\Sigma}}_{k}$ in (6) satisfies $(\epsilon_{0},0)$-DP given that
    $3\gamma_{k}\leq 2N_{k}\epsilon_{0}.$
    Proof:
    See [10, Theorem 4]. ∎
    Lemma 3 (Parallel composition).
    Suppose publishing $\widetilde{\boldsymbol{\mu}}_{k}$ and $\widetilde{\boldsymbol{\Sigma}}_{k}$ of an individual class $k$ satisfies $(\epsilon_{k}^{\prime},\delta)$-DP. Then, publishing $\{\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$ of all the classes satisfies $(\max_{k=1}^{K}\epsilon_{k}^{\prime},\delta)$-DP.
    Proof:
    The result directly follows from the standard parallel mechanism of DP; see, e.g., [5]. ∎
    Combining the above lemmas, it follows that releasing $\{\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$ adheres to $(\max_{k=1}^{K}\left\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\right\},\delta)$-DP.
    Next, we characterize the DP of releasing $\widetilde{\boldsymbol{\pi}}$. From Definition 1, it can be verified that $\widetilde{\boldsymbol{\pi}}$ satisfies $(\epsilon_{0},0)$-DP if the following holds for any adjacent dataset $D^{\prime}_{n,k^{\prime}}$:
    $e^{-\epsilon_{0}}\mathrm{Pr}(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}^{\prime}_{n,k^{\prime}}))\leq\mathrm{Pr}(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}))\leq e^{-\epsilon_{0}}\mathrm{Pr}(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}^{\prime}_{n,k^{\prime}})).$ (33)
    Combining the definition of $f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi})$, it follows that the condition in (33) is equivalent to (11).
    Finally, the overall DP can be obtained by using the following well-know composition theorem:
    Lemma 4 (Sequential composition).
    Let $M_{1}$ and $M_{2}$ be two randomized mechanisms that apply to an input dataset $\mathcal{D}$ and satisfy $(\epsilon_{1},\delta)$-DP and $(\epsilon_{2},0)$-DP, respectively. Then, the sequential execution of them satisfies $(\epsilon_{1}+\epsilon_{2},\delta)$-DP.
    Proof:
    See [5]. ∎
    According to Lemma 4, releasing $\{\tilde{\pi}_{k},\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k}\}_{k=1}^{K}$ satisfies $(\max_{k=1}^{K}\left\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\right\}+\epsilon_{0},\delta)$-DP.
    Appendix B Proof of Proposition 1
[…]
    $=$ $\mathbb{E}_{\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D})}\left[\sum_{k=1}^{K}\widetilde{p}(y=k)\left(\ln\frac{\widetilde{p}(y=k)}{p(y=k)}+\mathbb{E}_{\{\boldsymbol{\Gamma}_{k},\gamma_{k}\}_{\forall k}}\left[\int\widetilde{p}({\bf x}|y=k)\ln\frac{\widetilde{p}({\bf x}|y=k)}{p({\bf x}|y=k)}d{\bf x}\right]\right)\right]$
    $=$ $\sum_{\widetilde{\boldsymbol{\pi}}\in\mathcal{S}}f(\widetilde{\boldsymbol{\pi}}|\boldsymbol{\pi}(\mathcal{D}))\sum_{k=1}^{K}\tilde{\pi}_{k}\left(\ln\frac{\tilde{\pi}_{k}}{\pi_{k}}+\mathbb{E}_{\{{\bf w}_{k},{\bf W}_{k}\}_{\forall k}}\left[\int\mathcal{N}(\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k})\ln\frac{\mathcal{N}(\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k})}{\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})}d{\bf x}\right]\right).$ (34)
    We expend the KL divergence formula in (8) by substituting (1) and (7), yielding the expression in (34) shown on top of the next page. We note that the last term on the right-hand side of (34) is the KL divergence of two Gaussian distributions. Using the formula of the KL divergence between two multivariant Gaussian distributions [22], we have
    $\mathbb{E}\left[\int\mathcal{N}(\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k})\ln\frac{\mathcal{N}(\widetilde{\boldsymbol{\mu}}_{k},\widetilde{\boldsymbol{\Sigma}}_{k})}{\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})}d{\bf x}\right]$
    $=$ $\frac{1}{2}\mathbb{E}\left[{\bf w}_{k}^{T}\boldsymbol{\Sigma}_{k}^{-1}{\bf w}_{k}-d-\ln\frac{|\widetilde{\boldsymbol{\Sigma}}_{k}|}{|\boldsymbol{\Sigma}_{k}|}+\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1}\widetilde{\boldsymbol{\Sigma}}_{k})\right]$
    $=$ $\frac{1}{2}\left(\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1}\boldsymbol{\Gamma}_{k}^{-1})+\mathbb{E}\left[-\ln\frac{|\widetilde{\boldsymbol{\Sigma}}_{k}|}{|\boldsymbol{\Sigma}_{k}|}+\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1}{\bf W}_{k})\right]\right)$
    $=$ $\frac{1}{2}\left(\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1}\boldsymbol{\Gamma}_{k}^{-1})+\frac{d+1}{\gamma_{k}}\mathop{\rm tr}\nolimits(\boldsymbol{\Sigma}_{k}^{-1})+d\ln\gamma_{k}\right)$
    $-\frac{d\ln 2+\psi_{d}(\frac{d+1}{2})}{2},$ (35)
    where $\psi_{d}(\cdot)$ is the multivariant digamma function. Combining (34) and (35) completes the proof.
    Appendix C Proof of Proposition 2
    First, we prove by contradiction that the solution satisfies $F^{\prime}_{i,j^{\star}}>0$ for $\forall i$. Suppose there exist some $i^{\prime}$ such that $F^{\prime}_{i^{\prime},j^{\star}}=0$. The corresponding constraint becomes $e^{-\epsilon_{0}}F_{i,j}^{\prime}\leq F_{i,j^{\star}}=0$ implies $F_{i,j}^{\prime}=0$ for any $j\neq j^{\star}$. This contradicts with the row-stochastic constraint of $\sum_{j}F_{i,j}^{\prime}=1>0$.
    Next, define an auxiliary variable $\theta_{i,j}=\frac{F^{\prime}_{i,j}}{F^{\prime}_{i,j^{\star}}}$ for $\forall j\neq j$ and $\forall i$. The optimization problem is equivalent to:
    $\min_{\epsilon_{0},\{\theta_{i,j}\},\{F^{\prime}_{i,j^{\star}}\}_{\forall i}}\sum_{i}g_{i}F^{\prime}_{i,j^{\star}}$
    $\penalty 10000\ \text{s.t. }\penalty 10000\ \penalty 10000\ 0\leq\epsilon_{0}\leq\epsilon-\max_{k}\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\},$
    $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ e^{-\epsilon_{0}}\leq\theta_{i,j^{\star}}\leq e^{\epsilon_{0}},\forall j\neq j^{\star},\forall i.$
    $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ F_{i,j^{\star}}^{\prime}\left(1+\sum_{j\neq j^{\star}}\theta_{i,j}\right)=1,\forall i.$
    From the last two constraints, it follows that
    $F_{i,j^{\star}}^{\prime}=\frac{1}{1+\sum_{j\neq j^{\star}}\theta_{i,j}}\in\left[\frac{1}{1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{\epsilon_{0}}},\frac{1}{1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{-\epsilon_{0}}}\right].$
    Consequently, for any given $\epsilon_{0}$, the minimization of the linear objective over $\{F_{i,j^{\star}}\}_{\forall i}$ depends on each sign of $g_{i}$. Specifically, for any $i$ we consider the following two subcases:
    • $\bullet$
      If $g_{i}\geq 0$, the optimal solution is achieved when $F_{i,j^{\star}}^{\prime}=1/(1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{\epsilon_{0}})$.
    • $\bullet$
      If $g_{i}<0$, the optimal solution is achieved when $F_{i,j^{\star}}^{\prime}=1/(1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{-\epsilon_{0}})$.
    The final step is to find the optimal solution for $\epsilon_{0}$. Let $A=\sum_{i:g_{i}>0}g_{i}$ and $B=\sum_{i:g_{i}>0}g_{i}$. The problem can be recast as the minimization of a one-dimensional function $J(\epsilon_{0})=\frac{A}{1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{\epsilon_{0}}}+\frac{B}{1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{-\epsilon_{0}}}$ over the range of $0<\epsilon_{0}\leq\epsilon-\max_{k}\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\}$. The first-order derivative of $J(\epsilon_{0})$ is given by
    $J^{\prime}(\epsilon_{0})=\frac{Be^{\epsilon_{0}}|\mathcal{S}^{\prime}(\mathcal{D})|}{(1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{-\epsilon_{0}})^{2}}-\frac{Ae^{\epsilon_{0}}|\mathcal{S}^{\prime}(\mathcal{D})|}{(1+|\mathcal{S}^{\prime}(\mathcal{D})|e^{\epsilon_{0}})^{2}}\leq 0.$
    Consequently, $J(\epsilon_{0})$ is non-increasing in $\epsilon_{0}$, and the optimal solution is given by the upper bound of $\epsilon_{0}$.
    To summarize, the optimal solution is given by
    $\epsilon_{0}$ $=\epsilon-\max_{k}\{\epsilon_{k}+\frac{3\gamma_{k}}{2N_{k}}\},$
[…]
    $F_{i,j}^{\prime}$ $=\left\{\begin{aligned} e^{\epsilon_{0}}F_{i,j^{\star}}^{\prime},&\text{ if }g_{i}\geq 0\\ e^{-\epsilon_{0}}F_{i,j^{\star}}^{\prime},&\text{ otherwise}\\ \end{aligned}\right.,\forall j\neq j^{\star}.$